OpenStudy (anonymous):

Find exact value of cos (-5pi/12)

OpenStudy (imstuck):

You could use the identity cos (A + B)

OpenStudy (imstuck):

This would be cos(-45 + (-30)).

OpenStudy (imstuck):

That expands to\[\cos(-45)\cos(-30)-\sin(-45)\sin(-30)\]

OpenStudy (imstuck):

Each of those has values you can sub in.

OpenStudy (imstuck):

\[\cos(-45)=\frac{ \sqrt{2} }{ 2 }\]\[\sin(-45)=-\frac{ \sqrt{2} }{ 2 }\]\[\cos(-30)=\frac{ \sqrt{3} }{ 2 }\]\[\sin(-30)=-\frac{ 1 }{ 2 }\]

OpenStudy (imstuck):

So fill in your values now:

OpenStudy (imstuck):

\[(\frac{ \sqrt{2} }{ 2})(\frac{ \sqrt{3} }{ 2 })-(-\frac{ \sqrt{2} }{ 2 })(-\frac{ 1 }{ 2 })\]

OpenStudy (anonymous):

Thank you so much. I got \[\frac{ \sqrt{6 - \sqrt{2}} }{ 4 }\]

OpenStudy (imstuck):

That comes out to be\[\frac{ \sqrt{6}-\sqrt{2} }{ 4 }\]

OpenStudy (anonymous):

Thank you :)

OpenStudy (imstuck):

You're welcome!

OpenStudy (imstuck):

Crazy answer, but that is it!